# Lightning declines over shipping lanes following regulation of fuel sulfur emissions

### v4

@@ Introduction
Deep convective cloud (DCC) systems occur throughout the tropics, and are essential to the Earth’s water and energy cycles (Feng et al.,, 2021). However, there is little consensus on the mechanisms or magnitudes of aerosol particle impacts on DCCs (Tao et al.,, 2012; Seinfeld et al.,, 2016; Igel and van den Heever,, 2021; Varble et al.,, 2023). Thornton et al., (2017) documented a potential case of persistent maritime aerosol-DCC interactions analogous to stratocumulus ship tracks, with the discovery of enhancements in lightning over major shipping lanes passing through the Indian Ocean and South China Sea (Fig. 1).
@@
In January 2020, the International Maritime Organization (IMO) reduced the amount of allowable sulfur in fuel by a factor of seven, from 3.5% to 0.5% to curb effects of maritime shipping on air pollution (IMO,, 2020). Recent analyses of shallow stratocumulus marine clouds over shipping lanes find changes to cloud brightness, droplet number, and droplet size associated with the IMO regulation, presumably due to the shift in aerosol size distribution (Watson-Parris et al.,, 2022; Yuan et al.,, 2022; Diamond,, 2023).
@@
The Port of Singapore accounts for  20% of the world’s bunkering fuel demand. The two primary shipping lanes it services—the Indian Ocean and South China Sea (hereafter ”the shipping lanes”)—have nearly an order of magnitude higher traffic than other shipping lanes around the world (Figure 1, top panel) (Mao et al.,, 2022). As shown in Figure 1 (middle panel), prior to 2020, the mean absolute lightning stroke density measured by the World Wide Lightning Detection Network (WWLLN) remains enhanced over these shipping lanes, consistent with Thornton et al 2017. Since 2020, however, when the IMO regulation of sulfur emissions began, absolute lightning stroke density over the shipping lanes has declined by about 1 stroke km-2 year-1 as indicated by the difference between the means of years 2020-2023 and 2010-2019 (Figure 1, bottom panel). While some of the largest absolute declines in lightning since 2020 occur over the shipping lanes, lightning has increased or decreased in other parts of this region as well. As we illustrate below, variability in the dynamic and thermodynamic context for convection over these shipping lanes, as well as potential changes to shipping due to pandemic and other disruptions to trade must be taken into account to better isolate the potential impacts of shipping emissions.
Ship traffic in the vicinity, as measured by vessel fuel sales at the Port of Singapore, has been relatively constant or even increased since 2020 (Figure S1) (Port of Singapore,, 2024). The disruption by COVID-19 did not obviously decrease activity at the port, seeming only to have briefly slowed the growth of cargo throughput for 2-3 months in 2020 (Gu et al.,, 2023). Therefore, we focus on controlling for the variability in background meteorological conditions that impact the frequency and intensity of convection, and thus lightning, over the shipping lanes.
@@
We first examine the shipping lane lightning enhancement using two controls on background meteorology: 1) we only sample data from times and pixels with precipitating clouds(Huffman et al.,, 2015; Pradhan and Markonis,, 2023; Watters et al.,, 2023); and 2) we restrict analyses to the specific seasons in each region favorable for lighting (November to April in the Indian Ocean; June to November in the South China Sea). Using these criteria, we composite lightning observations as a function of distance to the shipping lanes, the center of which we define as the peak in shipping emissions from the EDGAR emissions inventory (see Methods). As shown in Figure 2, mean absolute lightning exhibits a clear enhancement over the shipping lane before 2020 (pre-IMO), between approximately 150km south to 150km north of the shipping lanes, and that has decreased since the regulation onset in 2020 (post-IMO) (Figure 2a). Consistent with Figure 1, lightning in this context has fallen region-wide, particularly to the south of the shipping lanes. We next examine how variability in convective activity between the two periods (pre and post IMO) has affected the regional lightning changes.
@@
To account for inter-annual variability in the frequency and intensity of convection in the region, we regress the observed annual lightning at a given distance from the shipping lane against three variables known to relate to lightning frequency: Convective Available Potential Energy (CAPE, discussed further below) and precipitation rates (Romps et al.,, 2018), as well as the annual mean Oceanic Niño Index (ONI). The influence of El Niño is assumed to be uniform across the domain for a given year. Inter-annual variability in the MJO was small and had a negligible impact when included in the regression (see SI). Further details on the regression are in Appendix A. The regressed variables explain  33% of variance of the annual means. We subtract the regressed lightning from the observed annual mean, leaving the anomalous mean lightning stroke density that cannot be explained by interannual variability in storm occurrence and intensity, shown in Figure 2b.
@@
The annual anomalous enhancement in lightning over the shipping lanes prior to 2020 is even more clear after regressing out meteorological variability, and so is the near step-change decrease in the anomaly after 2020 (2b). Prior to the IMO regulation, essentially 100% of fuel sold at the port was high-sulfur (2b, right axis); correspondingly, the lightning anomaly over the shipping lane was 4.0 strokes km-2 year-1 on average and was never below 3 strokes km-2 year-1 for more than one year at a time. Adoption of the IMO regulation was prompt in 2020, as indicated by the change in high-sulfur fuel from 100% to less than 35% of fuel sold at the Port of Singapore. The Port of Singapore experienced little attenuation of fuel sales at the onset of COVID-19 (Gu et al.,, 2023), and total fuel sales have increased since 2020 consistent with higher traffic (Figure S1). Since 2020, the shipping lane lightning enhancements compared to adjacent regions have declined by  50% to  2 strokes km-2 year-1 on average.
@@
Figure 2: (a) Lightning stroke density as a function of distance to the shipping lanes before and after the IMO regulation. Shading represents ±2SE and ±3SE (b) Hovmöller diagram of the annual mean lightning anomaly from the linear regression using Convective Available Potential Energy, precipitation, and Oceanic Niño Index from reanalysis data and observations (see text and SI for more details).
@@
To further control for higher-frequency variations in convective activity and intensity, we examine the lightning enhancement in a 2-dimensional CAPE and precipitation space, using 3-hourly coincident observations of CAPE, precipitation, and lightning. Romps et al., (2018) showed that CAPE $\times$ precipitation is a reasonable proxy for lightning frequency over land, and Cheng et al., (2021) showed that with a CAPE threshold, it is a reasonable proxy for tropical oceanic lightning frequency. The 3-hourly CAPE and precipitation observations implicitly capture variability arising from more indirect sources, such as sea surface temperatures (SST), MJO, fronts, etc (see SI for further discussion).
@@
We compute lightning frequency in each CAPE-Precipitation bin using data from a region centered over each shipping lane and from reference regions adjacent to the shipping lanes, similar to those in (Thornton et al.,, 2017) (see Figure S2). We then compute a relative enhancement in lightning over the shipping lanes, before and after the onset of the IMO regulation, by taking the difference between corresponding CAPE-Precipitation bin-means in the shipping lane and associated reference box. The resulting shipping lane lightning enhancements as a function of both CAPE and Precipitation are shown in Figure 3. Before the IMO regulation (Pre-IMO), a shipping lane lightning enhancement existed in nearly every thermodynamic setting (e.g., in each CAPE-Precipitation bin, Figure 3a,d) for both shipping lanes.
@@
Since the IMO regulation (Post-IMO), both shipping lanes exhibit significantly weaker lightning enhancements across most CAPE-Precipitation regimes (Figure 3b,e). The bin-by-bin differences between the Pre and Post-IMO lightning enhancement histograms are shown in Figure 3 (c and f). On average across all CAPE-Precip conditions, the lightning enhancement has decreased by 76% and by 47% for the Indian Ocean and South China Sea shipping lanes, respectively (Figure 3c,f). Thus, for the same convective setting characterized by CAPE and Precipitation rates, lightning stroke density over the shipping lanes is significantly less enhanced compared to adjacent regions after 2020 than it was prior to 2020.
@@
Based upon the above, we hypothesize that the decline in the lightning enhancement since 2020 is most consistent with the IMO regulation changing CCN in the region. If decreasing sulfur emissions over the shipping lanes has reduced the total number of viable CCN and disrupted an associated mechanism for lightning enhancement, then there should be a corresponding change in warm cloud microphysics. To further test our hypothesis, we use Moderate Resolution Imaging Spectroradiometer (MODIS) satellite observations of cloud droplet number (Nd) in warm (shallow) clouds over the Indian Ocean shipping lane, where the influence of land is weaker and ship emissions are stronger, during the high-lightning season. The retrievals of Nd follow the method outlined in Zhu et al., (2018), such that we sample only the brightest, most active warm clouds to ensure that assumptions of adiabatic ascent are as realistic as possible (see Appendix A).
Due to the optical thickness of DCC, the retrievals of Nd can only be done for shallow warm clouds. As a result, this analysis samples a different set of conditions than the lightning observations which arise essentially only from DCC. We thus assume that the behavior of Nd in shallow cumulus clouds from the same region is related to, though not necessarily a direct proxy for, Nd at cloud base in DCC. We have partially accounted for variability in aerosol particles from continental dust and biomass burning aerosol using MERRA-2 reanalysis estimates across the region in constructing Figure 4 to better isolate the effect of shipping emission changes on Nd. Limited observations in the region likely hinder the ability of reanalysis products to capture the full variability in CCN sources (see SI for discussion of aerosol optical depth), possibly explaining some remaining differences in Nd over the southern region of the domain pre and post 2020.
@@
In Figure 4, we show that prior to the IMO regulation, there is a clear trend in Nd toward land (north), as aerosol number concentrations are typically greater over land than over oceans (Figure S6), as well as a clear perturbation in Nd over the shipping lane. This Nd perturbation is roughly 10-15% above the average of droplet concentrations 150km north and 150km south, which is similar in magnitude to the shipping lane perturbations to Nd detected by Diamond et al., (2020) in Southeast Atlantic stratoculumus clouds. The Nd perturbation over the Indian Ocean shipping lane is a significant finding on its own, as observations of persistent, mean-state Nd perturbations by ships are rare (Diamond and Wood,, 2020), especially for convectively active regions we show here.
@@
Since the IMO regulation, the Nd away from the shipping lane mostly maintain their previous levels, as indicated by the overlap in the 95% confidence intervals (shading), particularly to the north. Meanwhile, the enhancement in Nd over the shipping lane has become essentially undetectable 4. The decline in Nd over the shipping lane relative to the surrounding region establishes additional support for the declining lightning enhancement being related to a shift in aerosol particle number-size distributions, and consequently changes to the CCN distributions over the shipping lanes induced by the IMO regulation. Nd derived from shallow cumulus clouds will not be directly proportional to the CCN available for activation in deep convective clouds (DCC) due to the different supersaturation conditions present in each cloud type ((Hobbs et al.,, 2000), nor to the lightning enhancements that might result from CCN enhancements. However, the change in Nd is an additional observable indication, independent of the lightning observations, that the IMO regulations have clearly shifted aerosol particle distributions over the shipping lanes of interest here.
@@
Figure 4: Warm cloud-base droplet number (Nd) concentrations over the Indian Ocean, derived from MODIS observations of optical depth and effective radius following the procedure from Zhu et al., (2018). The pre-IMO regulation period is 2010-2019. Note the general increasing trend northward through the domain toward greater land influence, and the broad localized enhancement over the shipping lane prior to the regulation. Shading represents ±2SE.
@@
The IMO regulation of ship fuel sulfur illustrates connections between international trade, air pollution, DCC microphysics, and lightning. Further work combining in situ and remote sensing of aerosol and cloud microphysics together with lightning frequency is needed to clarify the mechanisms behind these connections, and to quantify the relative roles of dynamic and microphysical responses. Our findings herein show that these regions remain a useful testbed for understanding aerosol pollution impacts on deep convective clouds and lightning.
@@
where X is the vector of predictors, CAPE, precipitation, and ONI, $\beta$ is the vector of coefficients. $\epsilon$ is the residual or ”anomaly” that is shown in the figure. This anomaly represents the difference between the lightning one would expect given the environmental conditions ($\beta*X$) (see Figure S3) and the observed lightning ($F$).
@@
where k represents the volume radius ratio of cloud droplets (rv) to re (k = (rv/re)³ = 0.8). The term fad denotes the adiabatic fraction, for which we assumed a constant value of 1 in our study, due to the absence of more refined alternatives (Bennartz and Rausch,, 2017; Grosvenor et al.,, 2018). Cw signifies the adiabatic cloud water condensation rate within an ascending cloud parcel, expressed in grams per cubic meter per meter (g m-3 m-1). The extinction efficiency factor, Qext, is assumed to be 2, and $\rho$w is the density of water. To enhance the accuracy of our Nd estimations for each 0.5° x 0.5° grid box, we excluded pixels where the solar zenith angle exceeded 65 degrees (Grosvenor and Wood, 2014). We also excluded of scenes containing mixed-phase, ice, or multilayer clouds. Consequently, after applying these filtering criteria, the remaining dataset comprised less than 1% of multilayer cloud pixels in any given grid. We use only the Indian Ocean shipping lane to maximize signal-to-noise, as the South China Sea has a much weaker signal due to its proximity to land and lower ship emissions. Inclusion of the South China Sea in the analysis does not alter the results. Finally, in order to remove the impact of dust storms advected over the Bay of Bengal, and to thereby reduce interannual variability in Nd outside the shipping lanes, we collocate 3-hourly MERRA-2 aerosol reanalysis output of dust and black carbon with the MODIS Nd retrievals. We then remove any Nd retrievals where dust concentrations increase above 1 ng m33{-3}- 3 or black carbon concentrations above 0.1 ng m33{-3}- 3 (approximately the 50th percentile in each case).
@@
ERA5 CAPE may be downloaded using the Copernicus API at cds.climate.copernicus.eu. IMERG Precipitation and MERRA-2 aerosol are available for download at disc.gsfc.nasa.gov. WWLLN data is available at WWLLN.net. ONI index is available at psl.noaa.gov/data/correlation/oni.data. Precipitation Feature reflectivity datasets are available for download at: https://atmos.tamucc.edu/trmm/data/. MODIS Aqua (MYD06) retrievals are available at ladsweb.modaps.eosdis.nasa.gov. Precipitation Feature reflectivity datasets are available for download at: https://atmos.tamucc.edu/trmm/data/. Global ship traffic density is available at: datacatalog.worldbank.org/search/dataset/0037580/Global-Shipping-Traffic-Density. Analysis and plotting available at 10.5281/zenodo.11373991 (Wright,, 2024).

### v5

@@ Introduction
Deep convective cloud (DCC) systems occur throughout the tropics, and are essential to the Earth’s water and energy cycles (Feng et al.,, 2021). However, there is little consensus on the mechanisms or magnitudes of aerosol particle impacts on DCCs (Tao et al.,, 2012; Seinfeld et al.,, 2016; Igel and van den Heever,, 2021; Varble et al.,, 2023; williams_contrasting_2002). Thornton et al., (2017) documented a potential case of persistent maritime aerosol-DCC interactions analogous to stratocumulus ship tracks, with the discovery of enhancements in lightning over major shipping lanes passing through the Indian Ocean and South China Sea (Fig. 1).
@@
In January 2020, the International Maritime Organization (IMO) reduced the amount of allowable sulfur in fuel by a factor of seven, from 3.5% to 0.5% to curb effects of maritime shipping on air pollution (IMO,, 2020). Recent analyses of shallow stratocumulus marine clouds over shipping lanes find changes to cloud brightness, droplet number, and droplet size associated with the IMO regulation, presumably due to the shift in aerosol number-size distribution (Watson-Parris et al.,, 2022; Yuan et al.,, 2022; Diamond,, 2023).
@@
The Port of Singapore accounts for  20% of the world’s bunkering fuel demand. The two primary shipping lanes it services—the Indian Ocean and South China Sea (hereafter ”the shipping lanes”)—have nearly an order of magnitude higher traffic than other shipping lanes around the world (Figure 1, top panel) (Mao et al.,, 2022). As shown in Figure 1 (middle panel), prior to 2020, the mean absolute lightning stroke density measured by the World Wide Lightning Detection Network (WWLLN) remains enhanced over these shipping lanes, consistent with Thornton et al., (2017). Since 2020, however, when the IMO regulation of sulfur emissions began, lightning over the shipping lanes has decreased to an annual stroke density about 1 stroke km-2 year-1 lower than before the regulation (Figure 1, bottom panel). While some of the largest absolute declines in lightning since 2020 occur over the shipping lanes, lightning has increased or decreased in other parts of this region as well. As we illustrate below, variability in the dynamic and thermodynamic context for convection over these shipping lanes must be taken into account to better isolate the potential impacts of shipping emissions.
Regional ship traffic, as measured by vessel fuel sales at the Port of Singapore, has been relatively constant or even increased since 2020 (Figure S1) (Port of Singapore,, 2024). The disruption by COVID-19 did not obviously decrease activity at the port, seeming only to have briefly slowed the growth of cargo throughput for 2-3 months in 2020 (Gu et al.,, 2023). Therefore, we focus on controlling for the variability in background meteorological conditions that impact the frequency and intensity of convection, and thus lightning, over the shipping lanes.
@@
We first examine the shipping lane lightning enhancement using two controls on background meteorology: 1) we only sample precipitating clouds(Huffman et al.,, 2015; Pradhan and Markonis,, 2023; Watters et al.,, 2023); and 2) we restrict analyses to the specific seasons in each region favorable for lighting (November to April in the Indian Ocean; June to November in the South China Sea). Using these criteria, we composite lightning observations as a function of distance to the shipping lanes, the center of which we define as the peak in shipping emissions from the EDGAR emissions inventory (see Methods). As shown in Figure 2a, mean absolute lightning exhibits a clear enhancement over the shipping lane before 2020 (pre-IMO), between approximately 150km south to 150km north of the shipping lanes, and that has decreased since the regulation onset in 2020 (post-IMO) (Figure 2a).
@@
To account for inter-annual variability in the frequency and intensity of convection in the region, we regress the observed annual lightning at a given distance from the shipping lane against three variables known to relate to lightning frequency (Convective Available Potential Energy (CAPE, discussed further below), precipitation rates (Romps et al.,, 2018), and the annual mean Oceanic Niño Index (ONI)) as well as several spatial variables such as latitude and longitude (Appendix A). Inter-annual variability in the MJO was small and had a negligible impact when included in the regression (see SI). The regressed variables explain 65% of variance of the annual means. We subtract the regressed lightning from the observed annual mean, leaving the anomalous mean lightning stroke density that cannot be explained by interannual variability in storm occurrence and intensity, shown in Figure 2b.
@@
The annual anomalous enhancement in lightning over the shipping lanes prior to 2020 is even clearer after regressing out meteorological variability, as is the near step-change decrease in the anomaly after 2020 (2b). Prior to the IMO regulation, essentially 100% of fuel sold at the port was high-sulfur (2b, right axis); correspondingly, the lightning anomaly over the shipping lane was 3.9 strokes km-2 year-1 on average and was never below 2.5 strokes km-2 year-1 for more than one year at a time. Adoption of the IMO regulation was prompt in 2020, as indicated by the change in high-sulfur fuel from 100% to less than 35% of fuel sold at the Port of Singapore. The Port of Singapore experienced little attenuation of fuel sales at the onset of COVID-19 (Gu et al.,, 2023), and total fuel sales have increased since 2020 consistent with higher traffic (Figure S1). Since 2020, the shipping lane lightning enhancements compared to adjacent regions have declined by  67% to 1.25 strokes km-2 year-1 on average.
@@
Figure 2: (a) Lightning stroke density composited as a function of distance to the shipping lanes before and after the IMO regulation. Shading represents ±2SE and ±3SE (b) Hovmöller diagram of the annual mean lightning anomaly from the linear regression using Convective Available Potential Energy, precipitation, and Oceanic Niño Index from reanalysis data and observations, as well as spatial variables (latitude, longitude, lat*lon, lat2, lon2)(see text and SI for more details).
@@
To further control for higher-frequency variations in convective activity and intensity, we examine the lightning enhancement in a 2-dimensional CAPE and precipitation space, using 3-hourly coincident observations of CAPE, precipitation, and lightning. Cheng et al., (2021), building on Romps et al., (2018), showed that CAPE×$\times$×precipitation is a reasonable proxy for tropical oceanic lightning frequency, given that a CAPE threshold is implemented. The 3-hourly CAPE and precipitation observations implicitly capture variability arising from more indirect sources, such as sea surface temperatures (SST), MJO events, fronts, etc (see SI for further discussion).
@@
We compute lightning frequency in each CAPE-Precipitation bin using data from a region centered over each shipping lane and from reference regions adjacent to the shipping lanes (see Figure S2). We then compute a relative enhancement in lightning over the shipping lanes, before and after the onset of the IMO regulation, by taking the difference between corresponding CAPE-Precipitation bin-means in the shipping lane and associated reference box. The resulting shipping lane lightning enhancements as a function of both CAPE and Precipitation are shown in Figure 3. Before the IMO regulation (Pre-IMO), a shipping lane lightning enhancement existed in nearly every thermodynamic setting (e.g., in each CAPE-Precipitation bin, Figure 3a,d) for both shipping lanes. As indicated by the larger pre-IMO perturbation, it seems that low-CAPE environments may have been more susceptible to aerosol enhancement of lightning.
@@
Since the IMO regulation (Post-IMO), both shipping lanes exhibit significantly weaker lightning enhancements across most CAPE-Precipitation regimes (Figure 3b,e). The bin-by-bin differences between the Pre and Post-IMO lightning enhancement histograms are shown in Figure 3 (c and f). On average across all CAPE-Precip conditions, the lightning enhancement has decreased by 76% and by 47% for the Indian Ocean and South China Sea shipping lanes, respectively (Figure 3c,f). That is, for the same convective setting characterized by CAPE and Precipitation rates, the enhancement in lightning over the shipping lanes (as compared to adjacent regions) is significantly smaller after 2020 than it was before 2020.
@@
Based upon the above, we hypothesize that the decline in the lightning enhancement since 2020 is most consistent with the IMO regulation changing CCN in the region. If decreasing sulfur emissions over the shipping lanes has reduced the total number of viable CCN and disrupted an associated mechanism for lightning enhancement, then there should be a corresponding change in warm cloud microphysics. To further test our hypothesis, we use Moderate Resolution Imaging Spectroradiometer (MODIS) satellite observations of cloud droplet number (Nd) in low clouds over the Indian Ocean shipping lane, where the influence of land is weaker and ship emissions are stronger, during the high-lightning season. The retrievals of Nd follow the method outlined in Zhu et al., (2018)(see Appendix A). Retrievals of Nd can only be done for shallow cumulus, not DCC. As a result, Nd retrievals sample a different set of conditions than the lightning observations. We assume that the behavior of Nd in shallow cumulus clouds from the same region is related to, though not necessarily a direct proxy for, Nd at cloud base in DCC.
@@
In Figure 4, we show that prior to the IMO regulation, there is a clear trend in Nd toward land (north), as well as a clear perturbation in Nd over the shipping lane. This Nd perturbation is roughly 10-15% above the average of droplet concentrations 150km north and 150km south, which is larger than the shipping lane perturbations to Nd detected by Diamond et al., (2020) in Southeast Atlantic stratoculumus clouds. The Nd perturbation over the Indian Ocean shipping lane is a significant finding on its own, as observations of persistent, mean-state Nd perturbations by ships are rare (Diamond and Wood,, 2020), especially for convectively active regions we show here.
@@
Since the IMO regulation, the Nd away from the shipping lane mostly maintain their previous levels, as indicated by the overlap in the 95% confidence intervals, particularly to the north (upwind). Meanwhile, the enhancement in Nd over the shipping lane has become essentially undetectable 4. The decline in Nd over the shipping lane relative to the surrounding region establishes additional support for a relationship between the declining lightning enhancement and a shift in aerosol particle number-size distributions over the shipping lanes induced by the IMO regulation. Nd derived from shallow cumulus clouds will not be directly proportional to the CCN available for activation in high-supersaturation DCC (Hobbs et al.,, 2000), nor to the lightning enhancements that might result from CCN enhancements. However, the change in Nd is an additional observable indication, independent of the lightning observations, that the IMO regulations have clearly shifted aerosol particle distributions over the shipping lanes of interest here.
@@
Figure 4: Warm cloud-base droplet number (Nd) concentrations over the Indian Ocean, derived from MODIS observations of optical depth and effective radius following the procedure from Zhu et al., (2018). The pre-IMO regulation period is 2010-2019. Note the general increasing trend northward through the domain toward greater land influence, and the broad localized enhancement over the shipping lane prior to the regulation. To remove the effects and trends of large biomass burning and dust events, we remove any Nd retrievals where dust concentrations increase above 1 µg m33{-3}- 3 or black carbon concentrations above 0.1 µg m33{-3}- 3 (approximately the 50th percentile in each case). Shading represents ±2SE.
@@
The IMO regulation of ship fuel sulfur illustrates connections between international trade, air pollution, DCC microphysics, and lightning. Further work combining in situ and remote sensing of aerosol and cloud microphysics together with lightning frequency is needed to clarify the mechanisms behind these connections, and to quantify the relative roles of dynamic and microphysical responses. Our findings herein show that these regions remain a useful testbed for understanding aerosol pollution impacts on DCC and lightning.
@@
where X is the vector of predictors, (CAPE, precipitation, ONI, latitude (lat), longitude (lon), lat*lon, lat2, lon2), $\beta$ is the vector of coefficients. $\epsilon$ is the residual or ”anomaly” that is shown in the figure. This anomaly represents the difference between the lightning one would expect given the environmental conditions ($\beta*X$) (see Figure S3) and the observed lightning ($F$). In accordance with Cheng et al., (2021) CAPE has been set to zero where CAPE1/2 ¡ 15 ms-1.
@@
where k represents the volume radius ratio of cloud droplets (rv) to re (k = (rv/re)³ = 0.8). The term fad denotes the adiabatic fraction, for which we assumed a constant value of 1 in our study, due to the absence of more refined alternatives (Bennartz and Rausch,, 2017; Grosvenor et al.,, 2018). Cw signifies the adiabatic cloud water condensation rate within an ascending cloud parcel, expressed in grams per cubic meter per meter (g m-3 m-1). The extinction efficiency factor, Qext, is assumed to be 2, and $\rho$w is the density of water. To enhance the accuracy of our Nd estimations for each 0.5° x 0.5° grid box, we excluded pixels where the solar zenith angle exceeded 65 degrees (grosvenor_effect_2014). We also excluded of scenes containing mixed-phase, ice, or multilayer clouds. Consequently, after applying these filtering criteria, the remaining dataset comprised less than 1% of multilayer cloud pixels in any given grid. We use only the Indian Ocean shipping lane to maximize signal-to-noise, as the South China Sea has a much weaker signal due to its proximity to land and lower ship emissions. Inclusion of the South China Sea in the analysis does not alter the results. Finally, in order to remove the impact of dust storms advected over the Bay of Bengal, and to thereby reduce interannual variability in Nd outside the shipping lanes, we collocate 3-hourly MERRA-2 aerosol reanalysis output of dust and black carbon with the MODIS Nd retrievals. We then remove any Nd retrievals where dust concentrations increase above 1 ng m33{-3}- 3 or black carbon concentrations above 0.1 ng m33{-3}- 3 (approximately the 50th percentile in each case). Limited observations in the region likely hinder the ability of reanalysis products to capture the full variability in CCN sources (see SI for discussion of aerosol optical depth), possibly explaining some remaining differences in Nd over the southern region of the domain pre and post 2020.
@@
ERA5 CAPE may be downloaded using the Copernicus API at cds.climate.copernicus.eu. IMERG Precipitation and MERRA-2 aerosol are available for download at disc.gsfc.nasa.gov. ONI index is available at psl.noaa.gov/data/correlation/oni.data. Precipitation Feature reflectivity datasets are available for download at: https://atmos.tamucc.edu/trmm/data/. MODIS Aqua (MYD06) retrievals are available at ladsweb.modaps.eosdis.nasa.gov. Precipitation Feature reflectivity datasets are available for download at: https://atmos.tamucc.edu/trmm/data/. Global ship traffic density is available at: datacatalog.worldbank.org/search/dataset/0037580/Global-Shipping-Traffic-Density. Analysis and plotting available at 10.5281/zenodo.11373991 (Wright,, 2024). WWLLN lightning location data are collected by a global scientific collaboration and managed by the University of Washington. The WWLLN collaboration receives no federal, state or private funds to pay for the network operations, which are fully paid for by data sales (available at https://wwlln.net). Therefore, the stroke-level data is not free to the public. The composited annual stroke densities (as a function of distance from the shipping lane) and the mean pre- and post-regulation stroke densities region-wide are provided as part of the Zenodo code supplement.